OpenStudy (richyw):

improper integral. \[\int^{\infty}_0 x^2e^{-x^2}\]

OpenStudy (richyw):

is there a quick method to show that this converges?

OpenStudy (anonymous):

limits. they both go "0:"

OpenStudy (richyw):

huh?

OpenStudy (anonymous):

\[\lim_{x\to0}f(x)=0\\{\rm and}\\ \lim_{x\to\infty}f(x)=0 \]

OpenStudy (anonymous):

and finite in between. so, the integral exists

OpenStudy (anonymous):

do you see it?

OpenStudy (richyw):

no. I don't.

OpenStudy (richyw):

let \(u=x^2\), \(du=2xdx\)\[\int^{\infty}_0\frac{ue^{-u}}{2\sqrt{u}}du\]\[\frac{1}{2}\int^{\infty}_0u^{1/2}e^{-u}du\]\[\frac{1}{2}\Gamma(3/2)\]

OpenStudy (richyw):

\[\frac{1}{2}\Gamma(1+1/2)=\frac{1}{2}\Gamma(1/2)=\frac{\sqrt{\pi}}{4}\]

OpenStudy (richyw):

I am so fed up with the section on improper integrals. the only way I can ever figure them out is by evaluating them. surely there must be a quick way to find a larger function that converges?

OpenStudy (agent0smith):

\[\LARGE \int\limits^{\infty}_0 \frac{x^2} {e^{x^2}}dx\] hmm maybe the comparison test http://tutorial.math.lamar.edu/Classes/CalcII/ImproperIntegralsCompTest.aspx

OpenStudy (richyw):

every resource I can find on the comparison test tells me the same thing. what I suck at is finding the function to compare!

OpenStudy (richyw):

ok here I think that \[\frac{1}{x^2}\geq \frac{1}{e^{x^2}} \quad \forall \; x\in [0,\infty )\]

OpenStudy (richyw):

and \[\int^{\infty}_0 \frac{1}{x^2}\text{d}x\]converges. So by comparison \[ \int\limits^{\infty}_0 \frac{x^2} {e^{x^2}}\text{d}x\]converges?

OpenStudy (richyw):

is this correct? is there a systematic way to find the comparison function for when taking an exam?

OpenStudy (anonymous):

Que 1) Does the function converge? \[\lim_{x\to0}x^2e^{-x^2}=0\\ \lim_{x\to\infty}x^2e^{-x^2}=0\times\lim_{x\to\infty}x^2=0\] So, the integral is finite.

OpenStudy (anonymous):

Que. 2) What is its value? \[ I={1\over2}\Gamma(1/2)=\Large \boxed{\pi\over4} \]

OpenStudy (richyw):

thanks. I really get worried because this seems to be straight out of calc 2! I have no idea what they are trying to teach me...

OpenStudy (anonymous):

i meant \[I=\Large{\sqrt{\pi}\over4}\]

OpenStudy (anonymous):

what do you mean? what calc class are you taking now?

OpenStudy (richyw):

uh it's called "Advanced Calculus and Introductory Analysis"

OpenStudy (richyw):

textbook is folland

OpenStudy (richyw):

it's basically revisiting the first three calculus courses from a different perspective and then doing some extra vector calculus stuff.

OpenStudy (anonymous):

I see. Then you should expect to see more of calc 1 & 2 combined. and calc 3 is nothing but the same and a little simpler. In this problem, what we have learnt is that the function might look non-converging but you cannot be sure unless you test the limits and for intermediate asymptotes. This looks like a differential of Gaussian distribution function

OpenStudy (agent0smith):

"In this problem, what we have learnt is that the function might look non-converging " It looks like a converging function to me, because the e^-x^2 exponent term approaches zero much faster than x^2, so the integral should converge to me \[\LARGE \int\limits\limits^{\infty}_0 \frac{x^2} {e^{x^2}}dx \]

OpenStudy (agent0smith):

x^2 approaches infinity much slower than e^x^2 on the bottom, so it should approach zero quickly and thus converge.